# Synthesizing Grid Data with Cyber Resilience and Privacy Guarantees

### v1

@@ I Introduction
Optimal power flow (OPF) analysis in power systems requires realistic grid models with accurate network, generation, and load parameters—data that is extremely difficult to source from real-world grids due to privacy and cybersecurity concerns. While the lack of such models has inspired the development of artificial grids [1, 2, 3], a more principled approach leverages the theory of differential privacy (DP) [4] to release grid models directly from real-world systems.
@@
However, the privacy guarantees alone may not suffice to release grid parameters, as cybersecurity risks associated with such releases remain largely unexplored. Possible cyber attacks include false data injection, which subtly alters state estimation results [14], and related load redistribution, which manipulates demand measurements to increase OPF cost and constraint violation [15]. While executing such attacks requires some grid knowledge [16], traditionally, this information has been difficult to obtain. However, the availability of synthetic grid parameters may unintentionally inform adversaries and help in calibrating the attack.
@@
Notation: lower- (upper-) case boldface letters denote column vectors (matrices). Scalar $a_{i}$ is the $i^{\text{th}}$ element of vector $\mathbf{a}$, while vector $\mathbf{a}_{i}$ collects the $i^{\text{th}}$ row of matrix $\mathbf{A}$. Vectors $\mathbf{0}$ and $\mathbf{1}$ are the all-zero and all-one vectors; $\top$ stands for transposition, and $\mathbf{x}^{\star}$ denotes the optimal value of $\mathbf{x}$.
@@ II Preliminaries
II-A Optimal Power Flow (OPF) Problem
The OPF problem seeks the least-cost generation dispatch that meets electrical loads while satisfying grid constraints:
@@ II-A Optimal Power Flow (OPF) Problem
subject to $\overline{\mathbf{p}}\leqslant\mathbf{p}\leqslant\overline{\mathbf{p}}$ (1b)
@@
The model maps the load vector $\mathbf{d}$ to the optimal dispatch cost $C_{\text{opf}}(\mathbf{d})$. The decision variables include the vector $\mathbf{p}$ of generator dispatch and vector $\mathbf{v}$ of power flow constraint violations. Dispatch costs are linear with coefficients $\mathbf{q}$. Constraint (1b) requires generator dispatch $\mathbf{p}$ to remain within a feasible range of $[\underline{\mathbf{p}},\overline{\mathbf{p}}]$. Constraint (1c) defines system-wide power balance between dispatched generation and loads. The power flows in transmission lines are computed using the matrix of power transfer distribution factors $\mathbf{F}$, which maps net power injections $(\mathbf{p}-\mathbf{d})$ to power flows as $\mathbf{F}(\mathbf{p}-\mathbf{d})$. They are capped by line capacity $\overline{\mathbf{f}}$ using constraint (1d). In highly loaded condition, these constraints can be temporally violated by $\mathbf{v}\geqslant\mathbf{0}$. Transmission constraint violations are not desired, so they are penalized in the objective function with a typically large parameter $\bm{\psi}$, i.e., $|\bm{\psi}|\gg|\mathbf{c}|$.
@@
where $\mathbf{x}=[\mathbf{p}^{\top}\;\mathbf{v}^{\top}]^{\top}$, $\mathbf{c}=[\mathbf{q}^{\top}\;\bm{\psi}^{\top}]^{\top}$. The $K$ inequalities in (2b) encode the dispatch constraints (1b), power flow constraints (1c) and (1d) using properly specified parameters $\mathbf{a}_{1},\mathbf{b}_{1},e_{1},\dots,\mathbf{a}_{K},\mathbf{b}_{K},e_{K}$. We make the following non-restrictive assumption on OPF.
Assumption 1.
For a feasible vector $\mathbf{d}$, problem (2) satisfies the linearly independent constraint qualification (LICQ).
@@ II-B Differential Privacy for Synthetic OPF Datasets
Optimization parameters in problem (1) are either classified or owned by private system actors, and thus can not be directly disclosed to public. Our goal is thus to synthesize some realistic version of these parameters. Without much loss of generality, we focus on the obfuscation of demand vector $\mathbf{d}$, while other parameters can be synthesized similarly. The state-of-the-art obfuscation algorithms synthesize high-quality demand vectors $\tilde{\mathbf{d}}$, which do not disclose the original vector $\mathbf{d}$ yet behave consistently in terms of power flow analysis [8, 6, 7]. They leverage DP to render original and synthetic vectors $\mathbf{d}$ and $\tilde{\mathbf{d}}$ statistically indistinguishable, up to some prescribed non-negative parameters $\alpha$ and $\varepsilon$. The former refers to adjacency of two arbitrary vectors [4].
@@
Two vectors $\mathbf{d},\mathbf{d}^{\prime}\in\mathcal{D}\subset\mathbb{R}^{n}$ are $\alpha-$adjacent, if $\exists i\in 1,\dots,n$, such that $d_{j}=d_{j}^{\prime},\forall j\in\{1,\dots,n\}\backslash i$, and $|d_{i}-d_{i}^{\prime}|\leqslant\alpha$ for $\alpha>0$. That is, they are different in one item by at most $\alpha$. $\triangleleft$
To synthesize a DP version $\tilde{\mathbf{d}}$ of $\mathbf{d}$, the standard Laplace mechanism applies a random noise to the original data, i.e., $\tilde{\mathbf{d}}=\mathbf{d}+\text{Lap}(\frac{\alpha}{\varepsilon})^{n}$, where $\text{Lap}(s)^{n}$ is the random draw from the $n-$dimensional Laplace distribution with zero mean and scale $s$ [7]. This mechanism satisfies the definition of $\varepsilon-$DP.
@@
A Laplacian mechanism is called $\varepsilon-$DP if, for any outcome $\tilde{\mathbf{d}}$ in the range of the mechanism, and any two $\alpha-$adjacent load vectors $\mathbf{d}$ and $\mathbf{d}^{\prime}$, the ratio of probabilities of observing the same outcome is bounded as
$\frac{\text{Pr}[\mathbf{d}{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}^{\prime}}+\text{Lap}(\frac{\alpha}{\varepsilon})^{n}=\tilde{\mathbf{d}}]}{\text{Pr}[\mathbf{d}^{\prime}+\text{Lap}(\frac{\alpha}{\varepsilon})^{n}=\tilde{\mathbf{d}}]}\leqslant\text{exp}(\varepsilon).$ (3)
@@
Parameter $\varepsilon$ is termed privacy loss. Intuitively, a smaller $\varepsilon$ results in more noise applied to data and higher requirement for distribution similarity, which would make it more likely to observe the same random outcome. However, the Laplace mechanism alone is likely to produce such load vector $\tilde{\mathbf{d}}$ that does not admit a feasible OPF solution, i.e., $C_{\text{opf}}(\tilde{\mathbf{d}})=\varnothing.$ The prior work introduced the following two-stage solution:
@@
$\underset{{\tilde{\mathbf{d}}}}{\text{minimize}}\;\;$ $\lVert C_{\text{opf}}(\tilde{\mathbf{d}})-\tilde{C}_{\text{opf}}\rVert_{1}+\gamma\lVert\tilde{\mathbf{d}}-\tilde{\mathbf{d}}^{0}\rVert_{1}$ (4a)
where the OPF costs come from
$C_{\text{opf}}(\tilde{\mathbf{d}})=\underset{{\mathbf{x}}}{\text{minimize}}\quad$ $\mathbf{c}^{\top}\mathbf{x}$ (4b)
subject to $\mathbf{a}_{k}^{\top}\mathbf{x}+\mathbf{b}_{k}^{\top}\tilde{\mathbf{d}}+e_{k}\leqslant\mathbf{0},\forall k$ (4c)
The synthetic vector $\tilde{\mathbf{d}}$ is optimized using feedback from the embedded OPF problem, which constraints $\tilde{\mathbf{d}}$ to take only those values that admit a feasible OPF solution. The main objective of (4) is to match the OPF cost on synthetic load vector with that on the original load vector, thereby ensuring high modeling fidelity of the synthetic data. For that purpose, (4) includes a differentially private estimate of OPF costs on true data $\tilde{C}_{\text{opf}}=C_{\text{opf}}(\mathbf{d})+\text{Lap}(\frac{2\overline{c}}{\varepsilon}),$ where $\overline{c}$ is the cost of the most expensive generator. The second term in (4a) is non-dominant and used to regularize the synthetic dataset with some small hyper-parameter $\gamma>0$. Solution to (4) is the feasible and cost-consistent synthetic counterpart $\tilde{\mathbf{d}}$, which ensures $\varepsilon-$DP guarantee for the original load vector $\mathbf{d}$ [8].
@@ III Cyber Resilience Risks in Releasing Differentially Private OPF Datasets
where the OPF costs come from
$C_{\text{opf}}(\mathbf{d}+\bm{\delta})=\underset{{\mathbf{x}}}{\text{minimize}}\;$ $\mathbf{c}^{\top}\mathbf{x}$ (5b)
subject to $\mathbf{a}_{k}^{\top}\mathbf{x}+\mathbf{b}_{k}^{\top}(\mathbf{d}+\bm{\delta})+e_{k}\leqslant\mathbf{0},\!\!\!$ (5c)
$\forall k.$ The attack vector $\bm{\delta}$ alters nominal loads using feedback from the embedded OPF problem, and is constrained by the set of admissible attacks
$\Delta$$\triangleq\left\{\bm{\delta}\;\bigg{|}\begin{array}[]{l}\underline{\bm{\delta}}\leqslant\bm{\delta}\leqslant\overline{\bm{\delta}}\\ \bm{1}^{\top}\bm{\delta}=0\end{array}\right\}$ (5f)
@@ IV Cyber Resilience and Privacy Guarantees for Synthetic OPF Datasets
$\underset{{\tilde{\mathbf{d}}}}{\text{minimize}}\;$ $\lVert C_{\text{opf}}(\tilde{\mathbf{d}})-\tilde{C}_{\text{opf}}\rVert_{1}+\beta\lVert C_{\text{att}}^{\text{BO}}(\tilde{\mathbf{d}})-\tilde{C}_{\text{opf}}\rVert_{1}$ (6)
where the first term control the fidelity of the synthetic data, and the second term measures the damage under attack, calibrated on the synthetic data. The objective is thus a trade-off between the fidelity of synthetic grid parameters and resilience of the grid to load redistribution attacks, which can be explored by varying parameter $\beta>0.$ The challenge, however, is that (6) requires solving a trilevel optimization problem, where the synthetic data is optimized over embedded BO attack model $C_{\text{att}}^{\text{BO}}(\tilde{\mathbf{d}})$. Inspired by [17], we seek computational tractability by exploring the connection between the bilevel model of attack and robust optimization.
@@ IV-A Computational Tractability via Robust Optimization (RO)
subject to $\underset{\bm{\delta\in\Delta}}{\max}\left[\mathbf{a}_{k}^{\top}\mathbf{x}+\mathbf{b}_{k}^{\top}(\mathbf{d}+\bm{\delta})+e_{k}\right]\leqslant\mathbf{0},\forall k,\!\!\!$ (7b)
where each constraint $k$ is formulated for the worst-case realization of the attack vector from the set of admissible attacks. In contrast to bilevel formulation (5), the RO attack is specific to each constraint. The following result shows that the RO attack provides an upper-bound on the BO attack.
@@
Although conservative, formulation (7) is computationally advantageous over (5) as it admits a linear programming reformulation via duality [18]. Let $\underline{\bm{\mu}}$ and $\overline{\bm{\mu}}$ be the duals of the first two constraints in (5f), and $\lambda$ be the dual of the last condition in (5f). The exact reformulation of (7) is
@@
Therefore, replacing $C_{\text{att}}^{\text{BO}}$ with $C_{\text{att}}^{\text{RO}}$ in objective function (6) gives rise to bilevel post-processing optimization, which can be handled by mixed-integer optimization solvers [8, 7].
@@ 3 Post-processing optimization of the synthetic load vector
The CRO algorithm for privacy-preserving and cyber-resilient synthesis of load parameters is summarized in Alg. 1. It takes as inputs load adjacency and $\varepsilon$-DP parameters, as well as optimization trade-off, regularization and attack parameters, $\beta$, $\gamma$ and $\Delta$, respectively. Step 1 initializes the synthetic load vector using the Laplace mechanism with a privacy loss of $\varepsilon_{1}$. Step 2 performs a DP estimation of the OPF costs on real loads using the Laplace mechanism with a privacy loss of $\varepsilon_{2}$. Following prior work in [8], this step requires the cost $\overline{c}$ of the most expensive generator. Finally, Step 3 post-processes the initial synthetic load by solving the bilevel optimization problem (1) using the conservative RO approximation of the attack. Since Step 3 does not optimize over real data, it does not introduce any privacy loss. The complete formulation of (1) can be seen in Appendix B.

### v2

@@ I Introduction
Optimal power flow (OPF) analysis in power systems requires realistic grid models with accurate network, generation, and load parameters—data that is difficult to source from real-world grids due to privacy and (cyber-)security concerns. While the lack of such models has inspired the development of artificial grids [1, 2], a more principled approach leverages the theory of differential privacy (DP) [3] to release grid models directly from real-world systems.
@@
However, the privacy guarantees alone may not suffice to release grid parameters, as cybersecurity risks associated with such releases remain largely unexplored. Possible cyber attacks include false data injection, which subtly alters state estimation results [13], line outage masking, which disconnects a transmission line and misguides a control center to seek outage elsewhere [14], and load redistribution, which manipulates demand measurements to increase OPF cost and constraint violation [15]. The latter is of main interest to this work. Executing such attacks requires some grid knowledge [16], which is traditionally difficult to obtain. However, the availability of synthetic grid data may unintentionally inform adversaries and help them calibrate the attack.
@@
Notation: lower- (upper-) case boldface letters denote column vectors (matrices). Scalar $a_{i}$ is the $i^{\text{th}}$ element of vector $\mathbf{a}$. Vectors $\mathbf{0}$ and $\mathbf{1}$ are the all-zero and all-one vectors; $\top$ stands for transposition, and $\mathbf{x}^{\star}$ is the optimal value of $\mathbf{x}$.
@@ II Preliminaries
II-A DC Optimal Power Flow (OPF) Problem
For a given load vector $\mathbf{d}$, the DC OPF problem seeks the least-cost generation dispatch in high-voltage grids that satisfies the loads and grid limits. Consider OPF as a parametric linear program:
@@ II-A DC Optimal Power Flow (OPF) Problem
subject to ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\underline{\mathbf{p}}}\leqslant\mathbf{p}\leqslant\overline{\mathbf{p}}$ (1b)
@@
where decision variables include generator dispatch $\mathbf{p}$, bounded by dispatch range $[\underline{\mathbf{p}},\overline{\mathbf{p}}]$, and power flow constraint violations $\mathbf{v}$, penalized by $\bm{\psi}$. The matrix $\mathbf{F}$ of power transfer distribution factors is used to map net power injections $(\mathbf{p}-\mathbf{d})$ to power flows as $\mathbf{F}(\mathbf{p}-\mathbf{d})$. Constraint (1c) defines system-wide power balance between dispatched generation and loads. The power flows in transmission lines are capped by line capacity $\overline{\mathbf{f}}$ using constraint (1d). In highly loaded condition, these constraints can be temporally violated by $\mathbf{v}\geqslant\mathbf{0}$. As transmission constraint violations are not desired, they are penalized with a large parameter $|\bm{\psi}|\gg|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{q}}|$.
@@
where $\mathbf{x}=[\mathbf{p}^{\top}\;\mathbf{v}^{\top}]^{\top}$, $\mathbf{c}=[\mathbf{q}^{\top}\;\bm{\psi}^{\top}]^{\top}$. The $K$ inequalities in (2b) encode the dispatch constraints (1b), power flow constraints (1c) and (1d) using properly specified parameters $\mathbf{a}_{1},\mathbf{b}_{1},e_{1},\dots,\mathbf{a}_{K},\mathbf{b}_{K},e_{K}$.
@@ II-B Differential Privacy for Synthetic OPF Datasets
Optimization parameters in problem (2) are either classified or owned by private system actors, and thus can not be directly disclosed to public. Our goal is thus to synthesize some realistic version of these parameters. In this work, we focus on the obfuscation of demand vector $\mathbf{d}$. This is without much loss of generality, because other parameters, such as transmission data in $\mathbf{a}_{k}$, $\mathbf{b}_{k}$ and $e_{k}$, can be synthesized similarly; see the state-of-the-art obfuscation algorithms [5, 6, 7]. Towards this goal, we leverage DP to render the original vector $\mathbf{d}$ statistically indistinguishable from its synthetic counterpart $\tilde{\mathbf{d}}$, up to some prescribed parameters: $\alpha$, termed the adjacency parameter, and $\varepsilon$, termed the privacy loss [3].
@@
Two vectors $\mathbf{d},\mathbf{d}^{\prime}\in\mathcal{D}\subset\mathbb{R}^{n}$ are $\alpha-$adjacent, for some $\alpha>0$, if $\exists i\in\{1,\dots,n$}, such that $d_{j}=d_{j}^{\prime},\forall j\in\{1,\dots,n\}\backslash i$, and $|d_{i}-d_{i}^{\prime}|\leqslant\alpha$. That is, they are different in one item by at most $\alpha$. $\triangleleft$
To synthesize a DP version $\tilde{\mathbf{d}}$ of $\mathbf{d}$, the standard Laplace mechanism applies a random noise to the original data, i.e., $\tilde{\mathbf{d}}=\mathbf{d}+\text{Lap}(\frac{\alpha}{\varepsilon})^{n}$, where $\text{Lap}(s)^{n}$ is a random draw from the $n-$dimensional Laplace distribution with zero mean and diagonal covariance matrix with each diagonal element equal $2s^{2}$ [6]. The mechanism guarantees that if the attacker’s prior for any load is within the ±$\pm\alpha$ MW range of the true value, it will not be improved by the DP release. If the prior is outside this range, the prior knowledge will be improved (thus enhancing grid transparency), but the exact loads will not be disclosed. In other words, the mechanism satisfies the following definition of $\varepsilon-$DP.
@@
The Laplace mechanism above, with domain $\mathcal{D}$ and output range $\mathcal{O}$, is called $\varepsilon-$DP if, for any outcome within $\hat{\mathcal{O}}\subseteq\mathcal{O}$ and any two $\alpha-$adjacent load vectors $\mathbf{d}$ and $\mathbf{d}^{\prime}$, the ratio of probabilities is bounded as
$\frac{\text{Pr}[\mathbf{d}{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}^{\prime}}+\text{Lap}(\frac{\alpha}{\varepsilon})^{n}\in\hat{\mathcal{O}}]}{\text{Pr}[\mathbf{d}^{\prime}+\text{Lap}(\frac{\alpha}{\varepsilon})^{n}\in\hat{\mathcal{O}}]}\leqslant\text{exp}(\varepsilon).$ (3)
@@
Intuitively, a smaller privacy loss $\varepsilon$ results in more noise applied to data and higher requirement for distribution similarity, which would make it more likely to observe the same random outcome. However, the Laplace mechanism alone is likely to produce such load vector $\tilde{\mathbf{d}}$ that does not admit a feasible OPF solution, i.e., $C_{\text{opf}}(\tilde{\mathbf{d}})=\varnothing.$ The prior work introduced the following two-stage solution:
@@
$\underset{{\tilde{\mathbf{d}}}}{\text{minimize}}\;\;$ $\lVert C_{\text{opf}}(\tilde{\mathbf{d}})-\tilde{C}_{\text{opf}}\rVert_{1}+\gamma\lVert\tilde{\mathbf{d}}-\tilde{\mathbf{d}}^{0}\rVert_{1}$ (4)
where the OPF costs $C_{\text{opf}}(\tilde{\mathbf{d}})$ comes from the embedded optimization problem (2) formulated on synthetic load vector $\tilde{\mathbf{d}}$, and $\tilde{C}_{\text{opf}}=C_{\text{opf}}(\mathbf{d})+\text{Lap}(\frac{2\overline{c}}{\varepsilon})$ computes a DP estimate of OPF costs on true data with $\overline{c}$ being the cost of the most expensive generator. The synthetic vector $\tilde{\mathbf{d}}$ is optimized using feedback from the embedded OPF problem, which constraints $\tilde{\mathbf{d}}$ to take only those values that admit a feasible OPF solution. The main objective of (4) is to match the OPF cost on synthetic load vector with that on the original load vector, thereby ensuring high modeling fidelity of the synthetic data. The second term in (4) is a regularization term with some small hyper-parameter $\gamma>0$ to choose the optimal solution that is closest to the original load after DP obfuscation $\tilde{\mathbf{d}^{0}}$. Solution to (4) is the feasible and cost-consistent synthetic counterpart $\tilde{\mathbf{d}}$, which ensures $\varepsilon-$DP guarantee for the original load vector $\mathbf{d}$ [7].
@@ III Cyber Resilience Risks in Releasing Differentially Private OPF Datasets
where the OPF costs $C_{\text{opf}}(\mathbf{d}+\bm{\delta})$ comes from the embedded optimization problem (2) formulated on load vector after attack $\mathbf{d}+\bm{\delta}$. The attack vector is constrained by the set of admissible attacks
$\Delta$$\triangleq\left\{\bm{\delta}\;\bigg{|}\begin{array}[]{l}\underline{\bm{\delta}}\leqslant\bm{\delta}\leqslant\overline{\bm{\delta}}\\ \bm{1}^{\top}\bm{\delta}=0\end{array}\right\}$ (5d)
@@ IV Cyber Resilience and Privacy Guarantees for Synthetic OPF Datasets
$\underset{{\tilde{\mathbf{d}}}}{\text{minimize}}\quad\lVert C_{\text{opf}}(\tilde{\mathbf{d}})-\tilde{C}_{\text{opf}}\rVert_{1}$ $+\beta\lVert C_{\text{att}}^{\text{BO}}(\tilde{\mathbf{d}})-\tilde{C}_{\text{opf}}\rVert_{1}$
$\quad{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+\gamma\lVert\tilde{\mathbf{d}}-\tilde{\mathbf{d}}^{0}\rVert_{1}}$ (6)
where the first term controls the fidelity of the synthetic data, the second term measures the damage under attack calibrated on the synthetic data, and the third term regularizes the demand vector. For a small penalty $\gamma$, this objective represents a trade-off between the fidelity of synthetic grid parameters and resilience of the grid to redistribution attacks, which can be explored by varying parameter $\beta>0.$ The embedded optimization $C_{\text{att}}^{\text{BO}}(\tilde{\mathbf{d}})$ includes the real grid data except for the load vector, thus modeling the worst-case attack when only the loads are unknown to adversaries.
The challenge is that (IV) requires solving a trilevel optimization problem, where the synthetic data is optimized over embedded BO attack model $C_{\text{att}}^{\text{BO}}(\tilde{\mathbf{d}})$. Inspired by [17], we seek computational tractability by exploring the connection between the bilevel model of attack and robust optimization.
@@ IV-A Computational Tractability via Robust Optimization (RO)
subject to $\underset{\bm{\delta_{k}\in\Delta}}{\max}\left[\mathbf{a}_{k}^{\top}\mathbf{x}+\mathbf{b}_{k}^{\top}(\mathbf{d}+\bm{\delta}_{k})+e_{k}\right]\leqslant\mathbf{0},\forall k,\!\!\!$ (7b)
where each constraint $k$ is formulated for the worst-case realization of the attack vector from the set of admissible attacks. In contrast to bilevel formulation (5), the RO attack generates a worst-case attack vector for each constraint. The following result shows that the RO attack provides an upper-bound on the BO attack.
@@
Although conservative, formulation (7) is computationally advantageous over (5) as it admits a linear programming reformulation via duality [18, §2.2] (see the link to online repository below for details). Let $\underline{\bm{\mu}}$ and $\overline{\bm{\mu}}$ be the duals of the first constraints in (5d), and $\lambda$ be the dual of the last condition in (5d). The exact reformulation of (7) is
@@
Therefore, replacing $C_{\text{att}}^{\text{BO}}$ with $C_{\text{att}}^{\text{RO}}$ in objective function (IV) gives rise to bilevel post-processing optimization, which can be handled by mixed-integer optimization solvers [7, 6].
@@ 3 Post-processing optimization of the synthetic load vector
The CRO algorithm for privacy-preserving and cyber-resilient synthesis of load parameters is summarized in Alg. 1. It takes as inputs load adjacency and $\varepsilon$-DP parameters, as well as optimization trade-off, regularization and attack parameters, $\beta$, $\gamma$ and $\Delta$, respectively. Step 1 initializes the synthetic load vector using the Laplace mechanism with a privacy loss of $\varepsilon_{1}$. Step 2 performs a DP estimation of the OPF costs on real loads using the Laplace mechanism with a privacy loss of $\varepsilon_{2}$. Following prior work in [7], this step requires the cost $\overline{c}$ of the most expensive generator. Finally, Step 3 post-processes the initial synthetic load by solving the bilevel optimization problem (1) using the conservative RO approximation of the attack. Since Step 3 does not optimize over real loads, it does not introduce any privacy loss. The complete formulation of (1) can be seen in Appendix B.